### v1

@@ 3.2.1 Multimodal Encoder
We discuss the feature extraction process of a multimodal encoder to create shared embedding space for both images and text. The multimodal encoder comprises two primary components: a text encoder and an image encoder. These components generate embeddings that capture the semantic relationships between images and corresponding textual descriptions. The image encoder is responsible for transforming input images into contextually enriched embeddings. On the other hand, the text encoder processes textual descriptions to produce corresponding embeddings. Utilizing transformer-based architectures, it captures contextual relationships within the text. The primary aim of the multimodal encoder is to align the embeddings of images and text within a shared semantic space. This alignment is achieved through a contrastive learning objective that encourages similar image and text pairs to have closer embeddings while pushing dissimilar pairs apart. We use a multimodal CLIP encoder [47] for this task.
@@ 3.2.2 Graph Knowledge Propagation
Graph-based knowledge propagation is the technique of disseminating information through a graph structure, where nodes signify features or data points, and edges depict relationships among them. It plays a pivotal role in tasks like information dissemination and graph-based learning. We present Algorithm 1, an approach for the construction of graphs and their knowledge propagation. The algorithm facilitates an iterative refinement of node representations of both image and text graph features. It encompasses a detailed procedure that includes the construction of graphs and the subsequent distribution of knowledge across various layers of the graph. In Algorithm 1, the operations from lines 1 to 19 are centered on constructing a similarity matrix ($A$). This process entails evaluating the cosine similarity between pairs of feature vectors denoted as $H_{i}$ and $H_{j}$. The decision to establish an edge between feature vectors depends on whether the computed cosine similarity $S_{ij}$ surpasses a predefined threshold (threshold $\geq$ 0.75). The resulting undirected edges are then incorporated into the adjacency matrix $A$. Consequently, lines 31 and 32 lead to the creation of two adjacency matrices: $A_{i}$, known as the Image similarity matrix, and $A_{t}$, known as the text similarity matrix. Subsequently, from lines 20 to 30, we employ GraphSAGE (Graph Sample and Aggregated) [22] for the propagation of knowledge across the constructed matrices. It is a node embedding algorithm designed specifically to acquire representations for nodes within a graph by aggregating information from their neighboring nodes. It strikes a balance between capturing local structural information and leveraging feature information associated with nodes. The fundamental concept behind this is to generate node embeddings by gathering information from the neighboring nodes of a given node. Here, each node represents a text feature in a text similarity matrix or an image in an image similarity matrix.
@@
27:Sample $N(v)$ from the neighborhood of node $v$ using the adjacency matrix $A$
@@
29:              $h_{v}^{k}\leftarrow\sigma\left(W_{k}\cdot\text{Concat}(h_{v}^{k-1},Agg^{k}_{N(v)})\right)$
@@
35:$A_{i}\leftarrow$ SimilarityGraph($H_{i}$, threshold)
36:$A_{t}\leftarrow$ SimilarityGraph($H_{t}$, threshold)
37:$I_{v}\leftarrow$ KnowledgeGraphPropagation($A_{i}$, $V_{i}$, $K$)
38:$T_{t}\leftarrow$ KnowledgeGraphPropagation($A_{t}$, $V_{t}$, $K$)
For a given node $v$ (representing a text or an image feature) and a specific layer $k$, the embedding generation process in GraphSAGE consists of the following steps, as depicted in lines 21-30: Neighbor Sampling, Aggregation, and Encoding. The process commences with the initialization of node embeddings $h_{0}^{v}$, instantiated from the similarity matrix ($A$). Subsequently, the algorithm systematically advances through $K$ layers, iteratively refining node representations. For each node $v$, a set of neighbors $N(v)$ is sampled from the adjacency matrix $A$. Information from the local neighborhood is then aggregated using a designated function ${Aggregate_{k}}$. The aggregation function can be as simple as using the mean to capture neighborhood information. This can be shown in Eq. 20.
@@
where, $h_{u}^{k-1}$ is the embedding of neighbor $u$ at layer $k-1$, and $\text{N}(v)$ is the set of sampled neighbors of node $v$. The updated node embeddings $h_{v}^{k}$ are computed by concatenating the aggregation result with the prior embeddings, facilitated by a trainable weight matrix $W_{k}$. The purpose of this step is to update the embedding of the central node by considering both its local neighborhood information and its own features. This can be represented as shown in Eq. 21:
$h_{v}^{k}\leftarrow\sigma\left(W_{k}\cdot\text{Concat}(h_{v}^{k-1},Agg^{k}_{N(v)})\right)$ (21)
@@ 3.2.3 Graph Fusion Learning Network (GFLN)
In the Graph Fusion Learning Network, the fusion of updated text and image graph features into a unified representation is achieved through concatenation. This is done to preserve discriminative information about each modality. The choice for this concatenation approach is rooted in the fact that we aim to obtain two types of features, i.e., cross-modal features and modality-specific refined features. Cross-modal features are obtained through the IDEA mechanism, enabling the model to capture and highlight inter-modal relationships. Concurrently, the graph learning module refines modality-specific features, emphasizing unique characteristics within each modality. Due to this strategy, the model ensures the preservation of both cross-modal interactions and modality-specific nuances, contributing to a comprehensive and enriched representation of the integrated information. Following the concatenation process, the unified vector achieves a dimensionality of 1024. Subsequently, we utilize three consecutive fully connected layers: the first layer reduces the dimensionality to 512 with a dropout rate of 0.2, followed by a second layer, further reducing it to 256 with the same dropout rate. The final layer further diminishes the dimensions to 128, incorporating a dropout layer set at 0.2. This processed vector is subsequently fed into the multimodal fusion module, enabling the seamless integration of information from diverse sources.
@@ 3.3.1.1 SentiQuotient Analysis (SentiQ):
SentiQ helps in deciphering the text’s subjective tone, thus contributing to a holistic interpretation of its underlying connotations. We use VADER (Valence Aware Dictionary and Sentiment Reasoner) [48] to extract the sentiments of the text. It assigns a sentiment score to each word in a piece of text. It computes a compound score that reflects the overall sentiment of the text, spanning from highly negative to highly positive, with neutrality positioned in the middle. It also provides scores for positive, negative, and neutral sentiments. By incorporating sentiment scores, models can better differentiate between classes by considering contextual sentiment and potentially weighing emotionally charged words more effectively. The SentiQ is represented as follows:
@@ 3.3.1.2 EmoQuotient Analysis (EmoQ):
EmoQuotient Analysis focuses on discerning emotional nuances expressed within the text, like happiness, sadness, anger, fear, etc. We use EmoLex [49] to quantitatively assess emotional content within textual data. Emotion Quotient plays a significant role in text classification by providing additional context and information about the emotional tone of the text. It gives 13 labels for a particular text. Due to this, we will get fine-grained scores for a text. These scores enhance the accuracy and granularity of the classification task. The EmoQ is represented as:
@@ 3.3.1.3 Crisis Informative Score (CIS):
$CIS=normalise(Crisis\_Lexicon\_Count)$ (24)
@@ 3.3.2.2 User Informative Score (UIS):
The User Informative Score is a metric designed to assess the credibility and informativeness of a user’s content. It is determined by considering the following factors: the total number of tweets posted by a user denoted as $U_{\text{Total}}$, the number of informative tweets posted by a user denoted as $U_{\text{Info}}$, and the number of non-informative tweets posted by a user, represented as $U_{\text{Non-Info}}$. As depicted in Algorithm 2, lines 1-5 tell about the computation of the UIS. It can be represented as shown in Eq. 25.
@@
The UIS metric is then normalized using min-max normalization. The Social Holistic Vector (SHV) comprises all the scores we got from Eq. 22,  Eq. 23,  Eq. 24, and  Eq. 25.
@@
As shown in Eq. 26, $\psi_{e},\psi_{s},\psi_{u},\psi_{u}i,\psi_{c}i$ represent the Emotion Quotient, Sentiment Quotient, User Engagement Metrics, User Informative Score, and Crisis Informative Score respectively.
@@ 3.4.1 Multimodal Adaptive Learning Network (MALN)
In this module, the fused attended vector (FAV) further refines and distills the integrated features by learning higher-level patterns and relationships that are specific to the classification task. The network can extract informative features and establish inter-dependencies between textual and visual features, ultimately enhancing the predictive capabilities of the model. The Fused Attended Vector (FAV), initially with dimensions of 1792, undergoes a hierarchical reduction in dimensionality. It passes through successive fully connected layers, achieving dimensions of 1024, 512, 256, and 128, respectively. Each layer incorporates a dropout rate of 0.2. The refined output is then directed to the Joint Fusion Learning Network for further processing.
@@ 3.4.2 Social Holistic Learning Network (SHLN)
The Social Holistic Vector (SHV) consolidates semantic text information and social interaction information, where the latter encompasses user engagement information and user-generated informative content shared on social media platforms. The user engagement data, including favorites count, retweet count, followers count, friends count, and status count, undergoes initial scaling using min-max normalization. This processed vector is referred to as the user engagement metrics. Similarly, user-generated informative content undergoes scaling through min-max normalization, yielding the User Informative Score (UIS), as outlined in subsection 3.3. The semantic textual information encompasses emotional, sentimental, and Crisis Informative Score (CIS). Following the concatenation of these processed feature vectors, the resulting comprehensive representation is denoted as the Social Holistic Vector, characterized by dimensions totaling 21. Specifically, it consists of 3 dimensions for sentiment, 11 for emotion, 1 for CIS, 5 for user engagement, and 1 for UIS. SHV encapsulates a holistic view of user interaction and content-related information on social media platforms. The SHV undergoes an initial processing stage by traversing through a fully connected layer, where its dimensions are effectively reduced to 16. To mitigate overfitting, a dropout of 0.2 is applied during this layer. Subsequently, another fully connected layer is employed, reducing the dimensions to 8 while maintaining the dropout rate at 0.2. These transformation includes dimensionality adjustments and the introduction of non-linearity through activation functions such as ReLU, enabling the network to capture intricate data patterns. During this process, the network extracts high-level features relevant to the classification task, including patterns related to user engagement, informative content, emotional and sentimental aspects. The resulting output represents a refined and condensed representation of the combined social and text information. This output is given to the Joint Fusion Learning Network.
@@ 3.4.3 Joint Fusion Learning Network (JFLN)
The feature vectors generated by MALN, SHLN, and GFLN are concatenated for joint representation. This integration enables the neural network to comprehensively capture the essence of multimodal data by merging diverse information streams. The joint representation encodes inter-modality relationships and interactions, allowing the model to discern complex patterns and correlations that may not be apparent when considering each modality in isolation. The joint representation of feature vectors passes through several fully connected layers; their dimensions undergo a step-wise reduction from 264 to 256, implementing a dropout of 0.2. This is succeeded by the application of another layer, further reducing the dimensions to 128 while maintaining a dropout rate of 0.2. The subsequent fully connected layer continues this reduction to 64 dimensions, maintaining the same dropout rate. This consolidated representation of the feature vectors is optimized for the classification task, empowering the model to make more accurate predictions by leveraging collective knowledge from all modalities, including the social context.
@@ 3.4.4 Prediction Layer
In the process of classifying content for the Informative Task, the model generates predicted class labels by passing the JFLN module’s output to the prediction layer. This layer employs a Sigmoid activation function to compute predicted probabilities, mapping real-valued inputs to probabilities between 0 and 1. Model performance is assessed using binary cross-entropy loss, and the Adam optimizer updates model weights based on loss gradients. The predicted probabilities are subsequently transformed into binary labels (0 or 1) using a thresholding function. If the probability surpasses or equals the threshold $\geq$ 0.5, it’s assigned the label 1, denoting the Informative class; otherwise, it’s labeled 0, indicating the non-informative class. This binary classification approach effectively distinguishes between informative and non-informative content.
@@
As shown in Eq. 27, $N$ represents number of examples, your dataset, $y_{i}$ represents true binary label (0 or 1) for the $i$-th example, $p_{i}$ represents predicted probability that the $i$-th example belongs to class 1.
@@
As shown in Eq. 28, $C$ represents number of classes, $N$ represents Number of examples, $y_{i}$ represents binary indicator (0 or 1) of whether class $i$ is the correct classification for the example, $p_{i}$ represents predicted probability assigned to class $i$ by the model, $y_{ij}$ represents binary indicator for class $i$ in the $j$-th example, $p_{ij}$ represents predicted probability for class $i$ in the $j$-th example.
@@
In summary, CrisisSpot is delineated into four integral components for the identification of crisis-related information. The primary Feature Extraction and Interaction Module orchestrates the interaction between text and visual modalities, extracted from text and image encoder, respectively, utilizing an IDEA attention mechanism. This mechanism adeptly captures both harmonious and contrary patterns, forwarding the extracted features to the subsequent Multimodal Fusion Module. The second component is dedicated to Multimodal Graph Learning, where features extracted from a multimodal encoder undergo enrichment of node features via a GraphSAGE layer. The resultant fused vector is channeled into the Multimodal Fusion Module. The third module is centered on Social Context Feature Extraction, encompassing diverse attributes such as crisis-related posts and user interaction history. This module includes metrics such as User Informative Score (UIS), Crisis Informative Score (CIS), and User Engagement. UIS evaluates user credibility, CIS quantifies crisis-related vocabulary presence, and User Engagement metrics capture pertinent text and user-centric information. The amalgamated features from this module are concatenated and fed into the Multimodal Fusion Module. The final module, Multimodal Fusion, processes the fused attended vector from the Feature Extraction Module and is passed through a dedicated Deep Neural Network, while a separate neural network handles the social holistic vector from the Social Context Feature Extraction Module. Outputs from these networks, in addition to graph-attended information, are further integrated through a Joint Fusion Learning Network, ultimately culminating in the prediction layer. This intricate architectural framework maximizes the synergistic utilization of textual, visual, and Social Context Features, augmenting the model’s prowess in discerning crisis-related information. CrisisSpot has a parameter size of 209.6M, and for processing each sample, it achieves a real-time inference speed of 4.25 microseconds.

### v2

@@ 4.2.1 Multimodal Encoder
We discuss the feature extraction process of a multimodal encoder to create shared embedding space for both images and text. The multimodal encoder comprises two primary components: a text encoder and an image encoder. These components generate embeddings that capture the semantic relationships between images and corresponding textual descriptions. The image encoder is responsible for transforming input images into contextually enriched embeddings. On the other hand, the text encoder processes textual descriptions to produce corresponding embeddings. Utilizing transformer-based architectures, it captures contextual relationships within the text. The primary aim of the multimodal encoder is to align the embeddings of images and text within a shared semantic space. This alignment is achieved through a contrastive learning objective that encourages similar image and text pairs to have closer embeddings while pushing dissimilar pairs apart. We use a multimodal CLIP encoder (Radford et al., 2021) for this task.
@@ 4.2.2 Graph Knowledge Propagation
Graph-based knowledge propagation is the technique of disseminating information through a graph structure, where nodes signify features or data points, and edges depict relationships among them. It plays a pivotal role in tasks like information dissemination and graph-based learning. We present Algorithm 1, an approach for the construction of graphs and their knowledge propagation. The algorithm facilitates an iterative refinement of node representations of both image and text graph features. It encompasses a detailed procedure that includes the construction of graphs and the subsequent distribution of knowledge across various layers of the graph. In Algorithm 1, the operations from lines 1 to 19 are centered on constructing a similarity matrix ($A$). This process entails evaluating the cosine similarity between pairs of feature vectors denoted as $H_{i}$ and $H_{j}$. The decision to establish an edge between feature vectors depends on whether the computed cosine similarity $S_{ij}$ surpasses a predefined threshold (threshold $\geq$ 0.75). The resulting undirected edges are then incorporated into the adjacency matrix $A$. Consequently, lines 31 and 32 lead to the creation of two adjacency matrices: $A_{i}$, known as the Image similarity matrix, and $A_{t}$, known as the text similarity matrix. Subsequently, from lines 20 to 30, we employ GraphSAGE (Graph Sample and Aggregated) (Hamilton et al., 2017) for the propagation of knowledge across the constructed matrices. It is a node embedding algorithm designed specifically to acquire representations for nodes within a graph by aggregating information from their neighboring nodes. It strikes a balance between capturing local structural information and leveraging feature information associated with nodes. The fundamental concept behind this is to generate node embeddings by gathering information from the neighboring nodes of a given node. Here, each node represents a text feature in a text similarity matrix or an image in an image similarity matrix.
@@
27:Sample $N(v)$ from the neighborhood of node $v$ using the adjacency matrix AA
@@
[…]
@@
[…]
For a given node $v$ (representing a text or an image feature) and a specific layer $k$, the embedding generation process in GraphSAGE consists of the following steps, as depicted in lines 21-30: Neighbor Sampling, Aggregation, and Encoding. The process commences with the initialization of node embeddings $h_{0}^{v}$, instantiated from the similarity matrix ($A$). Subsequently, the algorithm systematically advances through $K$ layers, iteratively refining node representations. For each node $v$, a set of neighbors $N(v)$ is sampled from the adjacency matrix $A$. Information from the local neighborhood is then aggregated using a designated function $Aggregate_{k}$}. The aggregation function can be as simple as using the mean to capture neighborhood information. This can be shown in Eq.˜21.
@@
where, $h_{u}^{k-1}$ is the embedding of neighbor $u$ at layer k1k-1, and $\text{N}(v)$ is the set of sampled neighbors of node $v$. The updated node embeddings $h_{v}^{k}$ are computed by concatenating the aggregation result with the prior embeddings, facilitated by a trainable weight matrix $W_{k}$. The purpose of this step is to update the embedding of the central node by considering both its local neighborhood information and its own features. This can be represented as shown in Eq.˜22:
[…]
@@ 4.2.3 Graph Fusion Learning Network (GFLN)
In the Graph Fusion Learning Network, the fusion of updated text and image graph features into a unified representation is achieved through concatenation. This is done to preserve discriminative information about each modality. The choice for this concatenation approach is rooted in the fact that we aim to obtain two types of features, i.e., cross-modal features and modality-specific refined features. Cross-modal features are obtained through the IDEA mechanism, enabling the model to capture and highlight inter-modal relationships. Concurrently, the graph learning module refines modality-specific features, emphasizing unique characteristics within each modality. Due to this strategy, the model ensures the preservation of both cross-modal interactions and modality-specific nuances, contributing to a comprehensive and enriched representation of the integrated information. Following the concatenation process, the unified vector achieves a dimensionality of 1024. Subsequently, we utilize two consecutive fully connected layers: the first layer reduces the dimensionality to 512 with a dropout rate of 0.2, followed by a second layer, further reducing it to 256 with the same dropout rate. This processed vector is subsequently fed into the multimodal fusion module, enabling the seamless integration of information from diverse sources.
@@ 4.3.1.1 SentiQuotient Analysis (SentiQ):
SentiQ helps in deciphering the text’s subjective tone, thus contributing to a holistic interpretation of its underlying connotations. We use VADER (Valence Aware Dictionary and Sentiment Reasoner) (Hutto and Gilbert, 2014) to extract the sentiments of the text. It assigns a sentiment score to each word in a piece of text. It computes a compound score that reflects the overall sentiment of the text, spanning from highly negative to highly positive, with neutrality positioned in the middle. It also provides scores for positive, negative, and neutral sentiments. By incorporating sentiment scores, models can better differentiate between classes by considering contextual sentiment and potentially weighing emotionally charged words more effectively. The SentiQ is represented as follows:
@@ 4.3.1.2 EmoQuotient Analysis (EmoQ):
EmoQuotient Analysis focuses on discerning emotional nuances expressed within the text, like happiness, sadness, anger, fear, etc. We use EmoLex (Mohammad and Turney, 2013) to quantitatively assess emotional content within textual data. Emotion Quotient plays a significant role in text classification by providing additional context and information about the emotional tone of the text. It gives 13 labels for a particular text (Anshul et al., 2023). Due to this, we will get fine-grained scores for a text. These scores enhance the accuracy and granularity of the classification task. The EmoQ is represented as:
@@ 4.3.1.3 Crisis Informative Score (CIS):
CIS=normalise(Crisis_Lexicon_Count)CIS=normalise(Crisis\_Lexicon\_Count) (25)
@@ 4.3.2.2 User Informative Score (UIS):
The User Informative Score is a metric designed to assess the credibility and informativeness of a user’s content. It is determined by considering the following factors: the total number of tweets posted by a user denoted as $U_{\text{Total}}$, the number of informative tweets posted by a user denoted as $U_{\text{Info}}$, and the number of non-informative tweets posted by a user, represented as $U_{\text{Non-Info}}$. As depicted in Algorithm 2, lines 1-5 tell about the computation of the UIS. It can be represented as shown in Eq.˜26.
@@
The UIS metric is then normalized using min-max normalization. The Social Holistic Vector (SHV) comprises all the scores we got from Eq.˜23, Eq.˜24, Eq.˜25, and Eq.˜26.
@@
As shown in Eq.˜27, $\psi_{e},\psi_{s},\psi_{u},\psi_{u}i,\psi_{c}i$ represent the Emotion Quotient, Sentiment Quotient, User Engagement Metrics, User Informative Score, and Crisis Informative Score respectively.
@@ 4.4.1 Multimodal Adaptive Learning Network (MALN)
In this module, the fused attended vector (FAV) further refines and distills the integrated features by learning higher-level patterns and relationships that are specific to the classification task. The network can extract informative features and establish inter-dependencies between textual and visual features, ultimately enhancing the predictive capabilities of the model. The Fused Attended Vector (FAV), initially with dimensions of 2048, undergoes a hierarchical reduction in dimensionality. It passes through successive fully connected layers, achieving dimensions of 1024, 512, and 256, respectively. Each layer incorporates a dropout rate of 0.2. The refined output is then directed to the Joint Fusion Learning Network for further processing.
@@ 4.4.2 Social Holistic Learning Network (SHLN)
The Social Holistic Vector (SHV) consolidates semantic text information and social interaction information, where the latter encompasses user engagement information and user-generated informative content shared on social media platforms. The user engagement data, including favorites count, retweet count, followers count, friends count, and status count, undergoes initial scaling using min-max normalization. This processed vector is referred to as the user engagement metrics. Similarly, user-generated informative content undergoes scaling through min-max normalization, yielding the User Informative Score (UIS), as outlined in subsection 4.3. The semantic textual information encompasses emotional, sentimental, and Crisis Informative Score (CIS). Following the concatenation of these processed feature vectors, the resulting comprehensive representation is denoted as the Social Holistic Vector, characterized by dimensions totaling 21. Specifically, it consists of 3 dimensions for sentiment, 11 for emotion, 1 for CIS, 5 for user engagement, and 1 for UIS. SHV encapsulates a holistic view of user interaction and content-related information on social media platforms. The SHV undergoes an initial processing stage by traversing through a fully connected layer, where its dimensions are effectively reduced to 16. To mitigate overfitting, a dropout of 0.2 is applied during this layer. These transformation includes dimensionality adjustments and the introduction of non-linearity through activation functions such as ReLU, enabling the network to capture intricate data patterns. During this process, the network extracts high-level features relevant to the classification task, including patterns related to user engagement, informative content, emotional and sentimental aspects. The resulting output represents a refined and condensed representation of the combined social and text information. This output is given to the Joint Fusion Learning Network.
@@ 4.4.3 Joint Fusion Learning Network (JFLN)
The feature vectors generated by MALN, SHLN, and GFLN are concatenated for joint representation. This integration enables the neural network to comprehensively capture the essence of multimodal data by merging diverse information streams. The joint representation encodes inter-modality relationships and interactions, allowing the model to discern complex patterns and correlations that may not be apparent when considering each modality in isolation. The joint representation of feature vectors passes through several fully connected layers; their dimensions undergo a step-wise reduction from 528 to 256, implementing a dropout of 0.2. This is succeeded by the application of another layer, further reducing the dimensions to 128 while maintaining a dropout rate of 0.2. The subsequent fully connected layer continues this reduction to 64 dimensions, maintaining the same dropout rate. This consolidated representation of the feature vectors is optimized for the classification task, empowering the model to make more accurate predictions by leveraging collective knowledge from all modalities, including the social context.
@@ 4.4.4 Prediction Layer
In the process of classifying content for the Informative Task, the model generates predicted class labels by passing the JFLN module’s output to the prediction layer. This layer employs a Sigmoid activation function to compute predicted probabilities, mapping real-valued inputs to probabilities between 0 and 1. Model performance is assessed using binary cross-entropy loss, and the AdamW optimizer updates model weights based on loss gradients. The predicted probabilities are subsequently transformed into binary labels (0 or 1) using a thresholding function. If the probability surpasses or equals the threshold $\geq$ 0.5, it’s assigned the label 1, denoting the Informative class; otherwise, it’s labeled 0, indicating the non-informative class. This binary classification approach effectively distinguishes between informative and non-informative content.
@@
As shown in Eq.˜28, $N$ represents number of examples, your dataset, $y_{i}$ represents true binary label (0 or 1) for the $i$-th example, $p_{i}$ represents predicted probability that the $i$-th example belongs to class 1.
@@
As shown in Eq.˜29, $C$ represents number of classes, $N$ represents Number of examples, $y_{i}$ represents binary indicator (0 or 1) of whether class $i$ is the correct classification for the example, $p_{i}$ represents predicted probability assigned to class $i$ by the model, $y_{ij}$ represents binary indicator for class $i$ in the $j$-th example, $p_{ij}$ represents predicted probability for class $i$ in the $j$-th example.
@@
In summary, CrisisSpot is delineated into four integral components for the identification of crisis-related information. The primary Feature Extraction and Interaction Module orchestrates the interaction between text and visual modalities, extracted from text and image encoder, respectively, utilizing an IDEA attention mechanism. This mechanism adeptly captures both harmonious and contrary patterns, forwarding the extracted features to the subsequent Multimodal Fusion Module. The second component is dedicated to Multimodal Graph Learning, where features extracted from a multimodal encoder undergo enrichment of node features via a GraphSAGE layer. The resultant fused vector is channeled into the Multimodal Fusion Module. The third module is centered on Social Context Feature Extraction, encompassing diverse attributes such as crisis-related posts and user interaction history. This module includes metrics such as User Informative Score (UIS), Crisis Informative Score (CIS), and User Engagement. UIS evaluates user credibility, CIS quantifies crisis-related vocabulary presence, and User Engagement metrics capture pertinent text and user-centric information. The amalgamated features from this module are concatenated and fed into the Multimodal Fusion Module. The final module, Multimodal Fusion, processes the fused attended vector from the Feature Extraction Module and is passed through a dedicated Deep Neural Network, while a separate neural network handles the social holistic vector from the Social Context Feature Extraction Module. Outputs from these networks, in addition to graph-attended information, are further integrated through a Joint Fusion Learning Network, ultimately culminating in the prediction layer. This intricate architectural framework maximizes the synergistic utilization of textual, visual, and Social Context Features, augmenting the model’s prowess in discerning crisis-related information.